# The Complexity of Resilience Problems via Valued Constraint Satisfaction Problems

### v3

@@ 8.2. Finite Duals
There is a finite dual ${\mathfrak{B}}$ of $\mu$ with domain $\{0,1\}$ which is finitely bounded homogeneous, as described in Example 8.2. That example also describes a valued structure $\Gamma$ which is actually $\Gamma({\mathfrak{B}},\emptyset)$. Let $\omega$ be the fractional cyclic operation given by $\omega(\min)=\omega(\max)=\frac{1}{2}$. Since $\operatorname{Aut}(\Gamma)$ is trivial, $\omega$ is canonical. The fractional operation $\omega$ improves both valued relations $R$ and $S$ (they are submodular; see, e.g., (Krokhin and Živný, 2017)) and hence is a canonical cyclic fractional polymorphism of $\Gamma$.
Combining Theorem 7.17 and 8.11, Example 8.13 reproves the results from (Freire et al., 2015b) (without multiplicities) and (Makhija and Gatterbauer, 2023) (with multiplicities) that the resilience problem for this query is in P.
@@ 8.3. Infinite Duals
The interesting direction is that if a finite $\sigma$-structure ${\mathfrak{A}}$ satisfies $\psi$, then ${\mathfrak{A}}$ embeds into $({\mathfrak{B}}_{\mu})_{\text{pp}(|\mu|)}$. Let $\phi$ be the canonical query of ${\mathfrak{A}}$. Let ${\mathfrak{A}}^{\prime}$ be the canonical database of the $\tau$-formula $\phi^{\prime}$. Suppose for contradiction that ${\mathfrak{A}}^{\prime}\models\mu$. Let $\chi$ be a minimal subformula of $\phi$ such that the canonical database of $\chi$ models $\mu$. Then $\chi$ has at most $|\mu|$ variables and implies $\mu$, and hence $\neg\chi$ is a conjunct of of $\psi$ which is not satisfied by ${\mathfrak{A}}$, a contradiction to our assumptions. Therefore, ${\mathfrak{A}}^{\prime}\models\neg\mu$ and by Property (2), we have that ${\mathfrak{A}}^{\prime}_{\text{pp}(|\mu|)}$ has an embedding $f$ into $({\mathfrak{B}}_{\mu})_{\text{pp}(|\mu|)}$.
@@
By Properties (1) and (6) of Theorem 8.14, ${\mathfrak{B}}_{\mu}$ is always a reduct of a finitely bounded homogeneous structure. For short, we write $\Gamma_{\mu}$ for $\Gamma({\mathfrak{B}}_{\mu},\emptyset)$ and $\Gamma_{\mu,\sigma}$ for $\Gamma({\mathfrak{B}}_{\mu},\sigma)$, see Definition 8.10. For some queries $\mu$, the structure ${\mathfrak{B}}_{\mu}$ can be replaced by a simpler structure ${\mathfrak{C}}_{\mu}$. This will be convenient for some examples that we consider later, because the structure ${\mathfrak{C}}_{\mu}$ is finitely bounded and homogeneous itself and hence admits the application of Theorem 7.17. To define the respective class of queries, we need the following definition. The Gaifman graph of a relational structure ${\mathfrak{A}}$ is the undirected graph with vertex set $A$ where $a,b\in A$ are adjacent if and only if $a\neq b$ and there exists a tuple in a relation of ${\mathfrak{A}}$ that contains both $a$ and $b$. The Gaifman graph of a conjunctive query is the Gaifman graph of the canonical database of that query.
@@
The resilience problem for a union of connected conjunctive queries $\mu$ where the relations from $\sigma\subseteq\tau$ are exogenous is polynomial-time equivalent to $\operatorname{VCSP}(\Gamma({\mathfrak{B}},\sigma))$ for any dual ${\mathfrak{B}}$ of $\mu$; in particular, to $\operatorname{VCSP}(\Gamma_{\mu,\sigma})$.
@@
To prove a polynomial-time reduction in the other direction, let $\phi$ be a $\tau$-expression. We construct a bag database ${\mathfrak{A}}$ with signature $\tau$. The domain of ${\mathfrak{A}}$ are the variables that appear in $\phi$ and for every $S\in\tau$ and tuple $(x_{1},\dots,x_{n})\in S^{{\mathfrak{A}}}$, its multiplicity in ${\mathfrak{A}}$ is the number of times that $(x_{1},\dots,x_{n})$ in $S^{{\mathfrak{A}}}$ occurs as a summand of $\phi$. The relations $S^{{\mathfrak{A}}}$ with $S\in\sigma$ are exogenous in ${\mathfrak{A}}$, the remaining ones are endogenous. Again, ${\mathfrak{A}}$ can be computed in polynomial time and the resilience of ${\mathfrak{A}}$ with respect to $\mu$ is at most $u$ if and only if $(\phi,u)$ has a solution over $\Gamma({\mathfrak{B}},\sigma)$. ∎
@@
We revisit Example 8.3. Consider the conjunctive query $\exists x,y,z\,(R(x,y)\wedge S(x,y,z))$ over the signature $\tau=\{R,S\}$. Note that the Gaifman graph of $\mu$ is complete; let ${\mathfrak{C}}_{\mu}$ be the structure from Theorem 8.15. We construct a binary pseudo cyclic canonical fractional polymorphism of $\Gamma({\mathfrak{C}}_{\mu},\emptyset)$. Let ${\mathfrak{M}}$ be the $\tau$-structure with domain $(C_{\mu})^{2}$ and where
@@
We proved that the corresponding valued structure has a binary canonical pseudo cyclic fractional polymorphism. By Theorem 7.17 and 8.16, this reproves the results from (Freire et al., 2015b) (without multiplicities) and (Makhija and Gatterbauer, 2023) (with multiplicities) that the resilience problem for this query is in P.

### v4

@@ 8.2. Finite Duals
There is a finite dual ${\mathfrak{B}}$ of $\mu$ with domain $\{0,1\}$ which is finitely bounded homogeneous, as described in Example 8.2. That example also describes a valued structure $\Gamma$ which is actually $\Gamma({\mathfrak{B}},\emptyset)$. Let $\omega$ be the fractional cyclic operation given by $\omega(\min)=\omega(\max)=\frac{1}{2}$. The fractional operation $\omega$ improves both valued relations $R$ and $S$ (they are submodular; see, e.g., (Krokhin and Živný, 2017)) and hence is a canonical cyclic fractional polymorphism of $\Gamma$. Combining Theorem 7.15 and 8.11, Example 8.12 reproves the results from (Freire et al., 2015b) (without multiplicities) and (Makhija and Gatterbauer, 2023) (with multiplicities) that the resilience problem for this query is in P.
In the following example we generalize the approach from Example 8.12.
Example 8.13.
Consider for $n\geq 1$ the conjunctive query
$\mu:=\exists x_{0},\dots,x_{n}(R_{1}(x_{0},x_{1})\wedge\cdots\wedge R_{n}(x_{n-1},x_{n})).$
Let $\tau=\{R_{1},\dots,R_{n}\}$ be the signature of $\mu$. We describe a finite dual ${\mathfrak{B}}$ of $\mu$. The domain $B$ consists of $2^{n-1}$ elements $v_{S}$, which we index by a subset $S$ of $[n-1]$. Then the dual satisfies $R_{i}(v_{S},v_{T})$ for $S,T\subseteq[n-1]$ and $i\in\{1,\dots,n\}$ if and only if
  • $i<n$ and $i\notin T$, or
  • $i>1$ and $(i-1)\in S$.
To see that ${\mathfrak{B}}$ is indeed a dual of $\mu$, let ${\mathfrak{A}}$ be a finite relational $\tau$-structure that does not satisfy $\mu$, and let $a\in A$. Let $S_{a}$ be the set of all elements $i\in[n-1]$ such that there exist elements $a_{i+1},\dots,a_{n}\in A$ such that
${\mathfrak{A}}\models R_{i+1}(a,a_{i+1})\wedge R_{i+2}(a_{i+1},a_{i+2})\wedge\cdots\wedge R_{n}(a_{n-1},a_{n}).$
We claim that $a\mapsto v_{S_{a}}$, $a\in A$, defines a homomorphism $h$ from ${\mathfrak{A}}$ to ${\mathfrak{B}}$.
Suppose that $a,b\in A$ are such that $R_{j}(h(a),h(b))$ does not hold in ${\mathfrak{B}}$ for some $j\in[n]$. This implies that
  • $j=n$ and $(n-1)\notin S_{a}$, or
  • $j=1$ and $1\in S_{b}$, or
  • $j\in\{2,\dots,n-1\}$, $(j-1)\notin S_{a}$, and $j\in S_{b}$.
In all three cases, the definition of $S_{a}$ and $S_{b}$ and ${\mathfrak{A}}\not\models\mu$ imply that $R_{j}(a,b)$ does not hold in ${\mathfrak{A}}$. This shows that $h$ is a homomorphism.
Now suppose that ${\mathfrak{A}}$ has a homomorphism to ${\mathfrak{B}}$. We have to show that ${\mathfrak{A}}$ does not satisfy $\mu$. It suffices to show that ${\mathfrak{B}}$ does not satisfy $\mu$. Suppose for contradiction that there are elements $v_{S_{0}},\dots,v_{S_{n}}$ such that ${\mathfrak{B}}$ satisfies $R_{1}(v_{S_{0}},v_{S_{1}})\wedge\cdots\wedge R_{n}(v_{S_{n-1}},v_{S_{n}})$. Then $R_{n}(v_{S_{n-1}},v_{S_{n}})$ implies that $(n-1)\in S_{n-1}$. By induction, we obtain from $R_{i}(v_{S_{i-1}},v_{S_{i}})$ that $(i-1)\in S_{i-1}$ for every $i\in\{2,\dots,n\}$. In particular, $1\in S_{1}$. However, note that $R_{1}(v_{S_{0}},v_{S_{1}})$ by definition implies that $1\notin S_{1}$. We reached a contradiction, hence, ${\mathfrak{B}}$ does not satisfy $\mu$.
Let $f$ be the symmetric binary operation that maps $(v_{S},v_{T})$ to $v_{S\cup T}$, and let $g$ be the symmetric binary operation that maps $(v_{S},v_{T})$ to $v_{S\cap T}$. Define the binary symmetric fractional operation $\omega$ by setting $\omega(f)=\omega(g)=\frac{1}{2}$. We claim that $\omega$ is a fractional polymorphism of $\Gamma({\mathfrak{B}},\emptyset)$.
Let $i\in\{1,\dots,n\}$. We have to show that for all $v_{S},v_{T},v_{P},v_{Q}\in B$, it holds in $\Gamma$ that
(13) $R_{i}(v_{S},v_{T})+R_{i}(v_{P},v_{Q})$ $\geq R_{i}(f(v_{S},v_{P}),f(v_{T},v_{Q}))+R_{i}(g(v_{S},v_{P}),g(v_{T},v_{Q}))$
$=R_{i}(v_{S\cup P},v_{T\cup Q})+R_{i}(v_{S\cap P},v_{T\cap Q}).$
First suppose that $1<i<n$. If the left-hand side in (13) is equal to $2$, then the disequality holds trivially. Whenever the left-hand side in (13) is equal to $1$, we must have $(i-1)\in S$, $(i-1)\in P$, $i\notin T$, or $i\notin Q$. Hence $(i-1)\in S\cup P$ or $i\notin T\cap Q$. In the first case, we have $R_{i}(v_{S\cup P},v_{T\cup Q})=0$, and in the second case we have $R_{i}(v_{S\cap P},v_{T\cap Q})=0$, and in both cases (13) holds. Finally, if the left-hand side in (13) is 00, then $R_{i}(v_{S},v_{T})=R_{i}(v_{P},v_{Q})=0$. Thus we have $(i-1)\in S$ or $i\notin T$ and at the same time $(i-1)\in P$ or $i\notin Q$. Therefore, we have $(i-1)\in S\cap P$, $i\notin T\cup Q$, or both $(i-1)\in S\cup P$ and $i\notin T\cap Q$. It follows that $R_{i}(v_{S\cup P},v_{T\cup Q})=R_{i}(v_{S\cap P},v_{T\cap Q})=0$. The case that $i=1$ and the case that $i=n$ can be treated similarly. Therefore, $\omega$ is a cyclic fractional polymorphism of $\Gamma({\mathfrak{B}},\emptyset)$, which by Theorem 8.11 implies that the resilience problem for $\mu$ is in P.
8.3. Regular Path Queries
We introduce two-way regular path queries and show that they have finite duals. A two-way regular path query (2RPQ) over a finite relational signature $\tau$ is a regular expression over the alphabet $\Sigma_{\tau}$ that consists of the symbols $R$ and $R^{-}$ for every binary relation symbol $R\in\tau$. We assume that regular expressions are built up from alphabet symbols, $\emptyset$, $\varepsilon$, $+$ (union), ; (concatenation), and $\cdot^{*}$ (Kleene star). As usual, we may omit the concatenation symbol for brevity.
Every 2RPQ $\mu$ defines a regular language $L(\mu)$ over the alphabet $\Sigma_{\tau}$. It also defines a database query of arity 2, as follows. A path in a $\tau$-database  $\mathfrak{A}$ is a sequence $a_{1}P_{1}a_{2}P_{2}\cdots P_{n-1}a_{n}$ with $a_{1},\dots,a_{n}\in A$ and $P_{1},\dots,P_{n-1}\in\Sigma_{\tau}$ such that for $1\leq i<n$, one of the following holds: (i) $P_{i}\in\tau$ and $\mathfrak{A}\models P_{i}(a_{i},a_{i+1})$ or (ii) $P_{i}=P^{-}$ with $P\in\tau$ and $\mathfrak{A}\models P(a_{i+1},a_{i})$. Note that we do not require $a_{1},\dots,a_{n}$ to be distinct. We call $a_{1}$ the source of the path, $a_{n}$ the target, and $P_{1}\cdots P_{n-1}$ the label. A pair $(a,b)\in A\times A$ is an answer to a 2RPQ $\mu$ if there is a path in $\mathfrak{A}$ that has source $a$, target $b$, and some label $w\in L(\mu)$. We use $\mu(\mathfrak{A})$ to denote the set of all answers to the 2RPQ $\mu$ on the database $\mathfrak{A}$.
We consider a Boolean version of the resilience problem for 2RPQs. For a 2RPQ $\mu$ over a signature $\tau$ and a set $\sigma\subseteq\tau$ of exogeneous relations, this problem is as follows: given a finite (bag) database ${\mathfrak{A}}$ and a $k\geq 1$, decide whether it suffices to remove at most $k$ tuples from ${\mathfrak{A}}$ to achieve that $\mu({\mathfrak{A}})=\emptyset$. It is folklore that given a 2RPQ $\mu$ and a database ${\mathfrak{A}}$, one can decide in polynomial time whether $\mu({\mathfrak{A}})=\emptyset$, e.g. by forming the product of ${\mathfrak{A}}$ and a finite automaton $A$ with $L(A)=L(\mu)$ and then checking whether some element with a final state in the second component is reachable from any element with the initial state in the second component. Consequently, for any 2RPQ $\mu$ with exogeneous relations from any set $\sigma$ the resilience problems is in NP by a straightforward guess-and-check algorithm.
Duals for 2RPQs are defined in exactly the same way as for conjunctive queries, that is, if $\mu$ is a 2RPQ with signature $\tau$, then a dual of $\mu$ is a $\tau$-structure ${\mathfrak{B}}$ such that any finite $\tau$-structure ${\mathfrak{A}}$ has a homomorphism to ${\mathfrak{B}}$ if and only if ${\mathfrak{A}}\not\models\mu$. We now argue that every 2RPQ has a finite dual. Note that the construction from (Calvanese et al., 2000) is related, but does not produce duals in our sense.
Lemma 8.14.
Every 2RPQ has a finite dual.
Proof.
Theorem 4.6 in (Bienvenu et al., 2014) states that every simple Boolean monadic Datalog program has a finite dual. It thus suffices to show that every 2RPQ can be expressed as such a program.
$\begin{array}[]{rcl}R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{n}(\bar{x}_{n})&\rightarrow&P(x)\\[2.84526pt] \textsf{true}(x)&\rightarrow&P(x)\\[2.84526pt] R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{n}(\bar{x}_{n})&\rightarrow&\textsf{goal}()\end{array}$
If $\mu$ is a 2RPQ and ${\mathfrak{B}}$ is the finite dual of $\mu$, then $\operatorname{VCSP}(\Gamma({\mathfrak{B}},\sigma))$ is polynomial-time equivalent to the Boolean resilience problem for $\mu$ with exogenous relations from $\sigma$. This can be proved in analogy with the proof of Theorem 8.11. The following result thus follows from the P versus NP-complete dichotomy theorem for valued VCSPs over finite domains stated in Theorem 7.15.
Theorem 8.15.
Let $\mu$ be a 2RPQ over a relational signature $\tau$ and let $\sigma\subseteq\tau$. Then the resilience problem for $\mu$ with exogenous relations from $\sigma$ is in P or NP-complete. Moreover, it is decidable whether the resilience problem for a given 2RPQ $\mu$ and set of exogeneous relations $\sigma$ is in P.
Example 8.16.
Consider the 2RPQ $\mu$ over the signature $\{R\}$ given by $R(RR^{-})^{*}RR$. It is well-known (see (Hell and Nešetřil, 2004)) that the relational $\{R\}$-structure ${\mathfrak{B}}$ with domain $\{0,1,2\}$ and the relation $R^{{\mathfrak{B}}}:=\{(0,1),(1,2)\}$ is a dual of $\mu$. Let $\Gamma:=\Gamma({\mathfrak{B}};\emptyset)$. Note that the crisp unary relation $\{0,1\}$ is expressible in $\Gamma$ by
$\phi(x):=\operatorname{Opt}(\inf_{y}R(x,y)).$
Let $R^{\prime}(x,y):=R(x,y)+\phi(x)+\phi(y)$ and note that $R^{\prime}\in\langle\Gamma\rangle$. It is easy to see that the valued structure ($(\{0,1,2\};R^{\prime})$ is homomorphically equivalent to the valued structure $\Gamma_{<}$ from Example 2.3. In other words, $\Gamma$ pp-constructs $\Gamma_{<}$. Since $\operatorname{VCSP}(\Gamma_{<})$ is NP-complete (Example 4.7), $\operatorname{VCSP}(\Gamma)$ is NP-complete by Lemma 5.13 (every finite-domain VCSP is in NP). Therefore, the resilience problem for the 2RPQ $\mu$ is NP-complete as well.
Example 8.17.
Consider the 2RPQ $\mu$ over the signature $\tau=\{R,S,T\}$ given by $RS^{*}T$. By Lemma 8.14, it has a finite dual. Consider the relational structure ${\mathfrak{B}}$ with domain $\{0,1\}$ and relations
  • $R^{{\mathfrak{B}}}=\{(0,0),(1,0)\}$,
  • $S^{{\mathfrak{B}}}=\{(0,0),(1,0),(1,1)\}$,
  • $T^{{\mathfrak{B}}}=\{(1,0),(1,1)\}$.
We argue that ${\mathfrak{B}}$ is a dual of $\mu$. Clearly, ${\mathfrak{B}}\not\models\mu$. Let ${\mathfrak{A}}$ be a finite $\tau$-structure with the domain $A$ such that ${\mathfrak{A}}\not\models\mu$. Let ${\mathfrak{A}}^{\prime}$ be the directed graph with the vertex set $A$ and the edge relation $E$ where $E=R^{\mathfrak{A}}\cup S^{{\mathfrak{A}}}\cup T^{{\mathfrak{A}}}$. We define $h\colon A\to\{0,1\}$ by setting $h(a)=0$ if there exists $b,c\in A$ such that $(b,c)\in R^{{\mathfrak{A}}}$ and there is a directed path from $c$ to $a$ in ${\mathfrak{A}}^{\prime}$, and $h(a)=1$ otherwise. It is straightforward to verify that $h$ is a homomorphism from ${\mathfrak{A}}$ to ${\mathfrak{B}}$.
Let $\omega$ be a binary fractional operation on $\{0,1\}$ defined by $\omega(\min)=\omega(\max)=1/2$, where $\min$ and $\max$ are the binary minimum and maximum operations on $\{0,1\}$. It is easy to check that $\omega\in\operatorname{fPol}(\Gamma({\mathfrak{B}},\emptyset))$. Therefore, $\omega$ is a cyclic fractional polymorphism of $\Gamma({\mathfrak{B}},\emptyset)$, which by Theorem 7.15 implies that $\operatorname{VCSP}(\Gamma({\mathfrak{B}},\emptyset))$ and therefore the resilience problem for $\mu$ is in P.
@@ 8.4. Infinite Duals
The interesting direction is that if a finite $\sigma$-structure ${\mathfrak{A}}$ satisfies $\psi$, then ${\mathfrak{A}}$ embeds into $({\mathfrak{B}}_{\mu})_{\text{pp}(|\mu|)}$. Let $\phi$ be the canonical query of ${\mathfrak{A}}$. Let ${\mathfrak{A}}^{\prime}$ be the canonical database of the $\tau$-formula $\phi^{\prime}$. Suppose for contradiction that ${\mathfrak{A}}^{\prime}\models\mu$. Let $\theta$ be a minimal subformula of $\phi^{\prime}$ such that the canonical database of $\theta$ models $\mu$. Then $\theta$ has at most $|\mu|$ variables and implies $\mu$, and hence $\neg\theta$ is a conjunct of of $\psi$ which is not satisfied by ${\mathfrak{A}}$, a contradiction to our assumptions. Therefore, ${\mathfrak{A}}^{\prime}\models\neg\mu$ and by Property (2), we have that ${\mathfrak{A}}^{\prime}_{\text{pp}(|\mu|)}$ has an embedding $f$ into $({\mathfrak{B}}_{\mu})_{\text{pp}(|\mu|)}$.
@@
By Properties (1) and (6) of Theorem 8.18, ${\mathfrak{B}}_{\mu}$ is always a reduct of a finitely bounded homogeneous structure. For short, we write $\Gamma_{\mu}$ for $\Gamma({\mathfrak{B}}_{\mu},\emptyset)$ and $\Gamma_{\mu,\sigma}$ for $\Gamma({\mathfrak{B}}_{\mu},\sigma)$, see Definition 8.10. For some queries $\mu$, the structure ${\mathfrak{B}}_{\mu}$ can be replaced by a simpler structure ${\mathfrak{C}}_{\mu}$. This will be convenient for some examples that we consider later, because the structure ${\mathfrak{C}}_{\mu}$ is finitely bounded and homogeneous itself. To define the respective class of queries, we need the following definition. The Gaifman graph of a relational structure ${\mathfrak{A}}$ is the undirected graph with vertex set $A$ where $a,b\in A$ are adjacent if and only if $a\neq b$ and there exists a tuple in a relation of ${\mathfrak{A}}$ that contains both $a$ and $b$. The Gaifman graph of a conjunctive query is the Gaifman graph of the canonical database of that query.
@@
The resilience problem for a union of connected conjunctive queries $\mu$ where the relations from $\sigma\subseteq\tau$ are exogenous is polynomial-time equivalent to $\operatorname{VCSP}(\Gamma({\mathfrak{B}},\sigma))$, for any dual ${\mathfrak{B}}$ of $\mu$; in particular, to $\operatorname{VCSP}(\Gamma_{\mu,\sigma})$.
@@
To prove a polynomial-time reduction in the other direction, let $\phi$ be a $\tau$-expression. We construct a bag database ${\mathfrak{A}}$ with signature $\tau$. The domain of ${\mathfrak{A}}$ are the variables that appear in $\phi$ and for every $S\in\tau$, we put a tuple $(x_{1},\dots,x_{n})$ in $S^{{\mathfrak{A}}}$ with a multiplicity equal to the number of occurences of $S(x_{1},\dots x_{n})$ as a summand of $\phi$. The relations $S^{{\mathfrak{A}}}$ with $S\in\sigma$ are exogenous in ${\mathfrak{A}}$, the remaining ones are endogenous. Again, ${\mathfrak{A}}$ can be computed in polynomial time and the resilience of ${\mathfrak{A}}$ with respect to $\mu$ is at most $u$ if and only if $(\phi,u)$ has a solution over $\Gamma({\mathfrak{B}},\sigma)$. ∎
@@
We revisit Example 8.3. Consider the conjunctive query $\exists x,y,z\,(R(x,y)\wedge S(x,y,z))$ over the signature $\tau=\{R,S\}$. Note that the Gaifman graph of $\mu$ is complete; let ${\mathfrak{C}}_{\mu}$ be the structure from Theorem 8.19 and recall that it is finitely bounded and homogeneous. We construct a binary fractional polymorphism of $\Gamma({\mathfrak{C}}_{\mu},\emptyset)$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}(\Gamma({\mathfrak{C}}_{\mu},\emptyset))=\operatorname{Aut}({\mathfrak{C}}_{\mu})$. Let ${\mathfrak{M}}$ be the $\tau$-structure with domain $(C_{\mu})^{2}$ and where
@@
We proved that the corresponding valued structure has a binary canonical pseudo cyclic fractional polymorphism. By Theorem 7.18 and 8.20, this reproves the results from (Freire et al., 2015b) (without multiplicities) and (Makhija and Gatterbauer, 2023) (with multiplicities) that the resilience problem for this query is in P.
We give another example of a conjunctive query with an infinite dual, which requires a ternary cyclic fractional polymorphism.
Example 8.23.
Consider the conjunctive query
$\mu:=\exists x,y\;(R(x,y)\wedge R(y,x))$
and observe that its resilience problem is in P: if ${\mathfrak{A}}$ is the input database, and both $(a,b)$ and $(b,a)$ lie in $R^{{\mathfrak{A}}}$ with multiplicities $m$ and $m^{\prime}$, respectively, we remove all copies of $(a,b)$ if $m\leq m^{\prime}$ and otherwise we remove all copies of $(b,a)$.
Let ${\mathfrak{C}}_{\mu}$ be the homogeneous dual of $\mu$ that embeds every countable structure ${\mathfrak{A}}$ that does not satisfy $\mu$. Let $C$ be the domain of ${\mathfrak{C}}_{\mu}$ and let $\Gamma:=\Gamma({\mathfrak{C}}_{\mu},\emptyset)$. We show that $\Gamma$ has a ternary canonical pseudo cyclic fractional polymorphism (with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$), which implies tractability of $\operatorname{VCSP}(\Gamma)$ (Theorem 7.18) and hence of the resilience problem for $\mu$ (Proposition 8.20). To increase readability, we write $R$ for $R^{{\mathfrak{C}}_{\mu}}$ and write $\breve{R}$ for $\{(a,b)\mid(b,a)\in R^{{\mathfrak{C}}_{\mu}}\}$. Note that since ${\mathfrak{C}}_{\mu}\models\neg\mu$, for every $a,b\in C$, we have precisely one of the following: $(a,b)\in R$, $(a,b)\in\breve{R}$, or $(a,b)\notin R\cup\breve{R}$. Let ${\mathfrak{M}}$ be an $\{R\}$-structure with the domain $C^{3}$ such that
${\mathfrak{M}}\models R((x,y,z),(u,v,w))$
if and only if at least one of the following is true:
  1. (1)
    at least two of $(x,u),(y,v),(z,w)$ lie in $R$;
  2. (2)
    $(x,u)\in R$, $(y,v)\in\breve{R}$, and $(z,w)\notin R\cup\breve{R}$;
  3. (3)
    $(x,u)\notin R\cup\breve{R}$, $(y,v)\in R$, and $(z,w)\in\breve{R}$;
  4. (4)
    $(x,u)\in\breve{R}$, $(y,v)\notin R\cup\breve{R}$, and $(z,w)\in R$.
Note that items (2)-(4) are just cyclic shifts of the same condition. It is straightforward to verify that ${\mathfrak{M}}\models\neg\mu$. For example, if ${\mathfrak{M}}\models R((x,y,z),(u,v,w))$ because of item (1), then at least two of $(u,x),(v,y),(w,z)$ lie in $\breve{R}$ and hence ${\mathfrak{M}}\models\neg R((u,v,w),(x,y,z))$ by definition. Therefore, there is an embedding $f$ of ${\mathfrak{M}}$ into ${\mathfrak{C}}_{\mu}$. By the definition of ${\mathfrak{M}}$, the operation $f$ is pseudo cyclic and canonical with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$ (this is easy to see, because ${\mathfrak{C}}_{\mu}$ is homogeneous; a detailed argument of this type can be found in Claim 6 in Proposition 8.27). The idea is that $f$ has the behavior of a majority operation on orbits of pairs of distinct elements.
Let ${\mathfrak{N}}$ be an $\{R\}$-structure with the domain $C^{3}$ such that
${\mathfrak{N}}\models R((x,y,z),(u,v,w))$
if and only if at least one of the following is true:
  1. (5)
    $(x,u),(y,v),(z,w)\in R$;
  2. (6)
    one of $(x,u),(y,v),(z,w)$ lies in $R$, and the remaining two lie in $\breve{R}$;
  3. (7)
    one of $(x,u),(y,v),(z,w)$ lies in $R$, and the remaining two do not lie in $R\cup\breve{R}$;
  4. (8)
    $(x,u)\in\tilde{R}$, $(y,v)\in R$, and $(z,w)\notin R\cup\breve{R}$;
  5. (9)
    $(x,u)\notin R\cup\breve{R}$, $(y,v)\in\breve{R}$ and $(z,w)\in R$;
  6. (10)
    $(x,u)\in R$, $(y,v)\notin R\cup\breve{R}$ and $(z,w)\in\breve{R}$.
Note that items (8)-(10) are just cyclic shifts of the same condition. Again one can verify that ${\mathfrak{N}}\models\neg\mu$, because ${\mathfrak{C}}_{\mu}\models\neg\mu$. Therefore, there is an embedding $g$ of ${\mathfrak{N}}$ into ${\mathfrak{C}}_{\mu}$. By the definition of ${\mathfrak{N}}$, the operation $g$ is pseudo cyclic and canonical with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$; it has the behavior of a minority operation on orbits of pairs of distinct elements.
Let $\omega$ be the ternary fractional operation defined by $\omega(f)=2/3$ and $\omega(g)=1/3$. Note that $\omega$ is a pseudo cyclic and canonical ternary fractional operation on $C$. We show that $\omega\in\operatorname{fPol}(\Gamma)$. Let $(x,u),(y,v),(z,w)\in C^{2}$. We want to verify that
$E_{\omega}\left[h\mapsto R\left(h\left(\begin{pmatrix}x\\ u\end{pmatrix},\begin{pmatrix}y\\ v\end{pmatrix},\begin{pmatrix}z\\ w\end{pmatrix}\right)\right)\right]\leq\frac{1}{3}(R(x,u)+R(y,v)+R(z,w)),$
equivalently,
(15) $2R\left(f\left(\begin{pmatrix}x\\ u\end{pmatrix},\begin{pmatrix}y\\ v\end{pmatrix},\begin{pmatrix}z\\ w\end{pmatrix}\right)\right)+R\left(g\left(\begin{pmatrix}x\\ u\end{pmatrix},\begin{pmatrix}y\\ v\end{pmatrix},\begin{pmatrix}z\\ w\end{pmatrix}\right)\right)\leq R(x,u)+R(y,v)+R(z,w).$
We break into cases:
  • If $(x,u),(y,v),(z,w)\in R$, then by item (1) and (5) the left-hand side of (15) evaluates to 0 and hence (15) holds.
  • If exactly two of $(x,y),(y,v),(z,w)$ lie in $R$, then by item (1) the left-hand side of (15) is at most 1 and hence (15) holds.
  • If exactly one of $(x,y),(y,v),(z,w)$ lies in $R$, then precisely one of the conditions (2)-(4), (6)-(10) applies and therefore the left-hand side of (15) is at most 2. Therefore, (15) holds.
  • If $(x,u),(y,v),(z,w)\notin R$, then (15) holds trivially since the left-hand side is always at most 3.
We conclude that $\omega\in\operatorname{fPol}(\Gamma)$.
We remark that there is no binary fractional polymorphism of $\Gamma$ that is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$; this is shown in detail in (Žaneta Semanišinová, 2025, Example 5.23).